OpenStudy (anonymous):

What is the solution to the equation 21 ^x – 3 = 14 ?

OpenStudy (hoblos):

21^x = 11 ln(21)^x = ln11 xln21= ln11 x= (ln11)/(ln21)

OpenStudy (anonymous):

which one? x ≈ –2.133 x ≈ –1.846 x ≈ 3.867 x ≈ 4.154

OpenStudy (hoblos):

ooppsss sorry sorry it is 21^x = 17 ln(21)^x = ln17 xln21= ln17 x= (ln17)/(ln21)

OpenStudy (anonymous):

Thank you! Appreciate it

OpenStudy (hoblos):

welcome :D

OpenStudy (anonymous):

i think its B